Given the following market​ equations: ​

Supply: Qs ​ = -12   ​+ 3 p

Demand: Qd=88-2p

Solve for the equilibrium price​ = ​$?? .   ​(round your calculation to the nearest penny​)

Concept Question 3.3 Question Help Given the following market equations: Supply: Qs123p Demand: Q 88 -2p Solve for the equilibrium price-$. (round your calculation to the nearest penny)
Answer #1

Equilibrium price can be determined at a point where

Supply = Demand

Given, Qs = - 12+ 3p

Qd = 88- 2p

Qs = Qd

- 12 + 3p = 88 - 2p

5p = 100

p = 100

p = $20

Therefore, equilibrium price is $20.